# Secure Smart Contract with Control Flow Integrity

### v1

@@ 6. Evaluation
In this section, we conduct a detailed assessment of CrossGuard. Our evaluation aims to answer the following research questions:
@@ 2. RQ 3:
    How do various actors, aside from hackers, introduce new control flows?
@@ 5. RQ 6:
  bZx Lending (bZx, 2025) oracle manipulation 7
29037 49 1707 24 16061 28 7736 18 3533 14
Warp Lending (War, 2025) oracle manipulation 17     ✓     416 1 11 0 404 0 0 0 1 1
CheeseBank Lending (Che, 2025) oracle manipulation 12     ✓     2404 1 125 0 1690 0 543 0 46 1
InverseFi Lending (Inv, 2025) oracle manipulation 12     ✓     126652 73 731 12 51471 17 46680 2 27770 44
CreamFi1 Lending (Cre, 2025a) re-entrancy 6     ✓     190865 121 321 32 181677 25 5795 34 3072 55
CreamFi2 Lending (Cre, 2025b) oracle manipulation 24     ✓     220962 273 500 56 204485 40 5073 33 10904 166
RariCapital1 Lending (Rar, 2025a) read-only re-entrancy 8     ✓     7142 13 58 1 6831 5 209 10 44 5
RariCapital2 Lending (Rar, 2025b) re-entrancy 12     ✓     84485 27 424 2 45217 0 27700 10 11144 19
XCarnival Yield Earning (XCa, 2025) logic flaw 6     ✓     877 3 61 0 800 0 0 0 16 3
Harvest Yield Earning (Har, 2025) oracle manipulation 11     ✓     31002 9 623 5 30338 2 2 0 39 2
ValueDeFi Yield Earning (Val, 2025) oracle manipulation 9     ✓     362 1 99 0 262 0 0 0 1 1
Yearn Yield Earning (Yea, 2024) logic flaw 7     ✓     133704 38 2855 10 81023 18 45035 8 4791 15
VisorFi Yield Earning (Vis, 2025) re-entrancy 3
86777 0 55 0 26480 0 47314 0 12928 0
UmbrellaNetwork Yield Earning (Umb, 2025) integer underflow 1     ✗     61 0 3 0 57 0 0 0 1 0
PickleFi Yield Earning (Pic, 2025) fake tokens 4     ✓     7830 2 1402 0 5168 0 1231 1 29 1
Eminence DeFi (Emi, 2025) logic flaw 6     ✓     22542 1 25 0 9473 0 9968 0 3076 1
Opyn DeFi (Opy, 2025) logic flaw 4
3937 1 29 1 607 0 1 0 3300 0
IndexFi DeFi (Ind, 2025) logic flaw 6     ✓     70735 5 98 0 14981 4 28186 0 27470 1
RevestFi Yield Earning (Rev, 2025) re-entrancy 5     ✓     2186 6 32 0 2127 5 19 0 8 1
DODO DeFi (DOD, 2025b) access control 2
1523 1 2 1 1285 0 195 0 41 0
Punk NFT (Pun, 2025) access control 3     ✓     111 4 14 1 96 2 0 0 1 1
BeanstalkFarms DAO (Bea, 2025) DAO attack 8     ✓     58648 13 243 9 43713 1 8923 0 5769 4
DoughFina Lending (Dou, 2025) no input validation 2     ✓     19 0 16 0 2 0 0 0 1 0
Bedrock_DeFi Restaking (Bed, 2025) price miscalculation 3     ✗     3426 4 3 0 2127 0 893 0 403 4
OnyxDAO DAO (Ony, 2025) fake market 8     ✓     157442 7 225 2 97210 0 393 0 59614 5
BlueberryProtocol Yield Earning (Blu, 2025) decimal difference 5     ✓     493 2 124 1 365 0 0 0 4 1
PrismaFi Restaking (Pri, 2025) no input validation 3     ✓     43669 34 135 2 24722 22 4187 0 14625 24
PikeFinance Lending (Pik, 2025) uninitialized proxy 1     ✓     8411 1 18 0 7026 0 0 0 1367 1
GFOX Game Fi (GFO, 2025) access control 2     ✓     12442 1 30 0 9392 0 12 0 3008 1
UwULend Lending (UwU, 2025) oracle manipulation 1     ✓     19765 101 178 3 18298 43 7 3 1282 84
                                  Avg. Ratio 100 100 6.81 19.7 68.34 18.4 12.67 8.92 12.18 52.78
@@ 6.1. Methodology
Hacked Protocol Selection: To address the research questions, we selected hacked DeFi protocols experiencing significant financial losses (exceeding \$300k) on Ethereum. Our dataset comprises two sets. The first set of 21 distinct victim protocols  [^6] Harvest Finance experienced two separate hacks on different contracts but is considered one protocol for our analysis. were identified by (Chen et al., 2024) from February 14, 2020, to August 1, 2022, excluding four cross-chain bridge hacks as they fall outside our scope. The second set comprises $8$ hacked protocols identified between February and July 2024 from DeFiHackLabs (Contributors, 2025). Table 2 details the selected benchmarks, presenting their Protocol Type, the hack transaction (Hack), the exploited vulnerability (Hack Type), and the number of affected contracts (#C). These vulnerabilities include diverse issues such as oracle manipulation, re-entrancy, integer underflows, access controls, and incorrect input validation. The selected protocols represent a broad spectrum of DeFi categories.
Target Contract Selection and Transaction History Retrieval: After identifying the hack transactions for each victim protocol, we collected all associated contracts involved. We used labels from Etherscan (Eth, 2024) to accurately determine the contracts belonging to the victim protocol by examining deployer addresses. We then retrieved the complete transaction history of these identified contracts from their deployment until the hack event. This comprehensive transaction history, including the hack transaction itself, constitutes the operational dataset of each victim protocol.  [^7] Although we might miss a few affected protocol contracts due to varying deployers or indirect involvement, having more contracts to protect would only introduce more complexity to the control flows, not reduce it. Thus, our analysis remains valid even with a subset of core contracts.
Control Flow Extraction and Simplification: To analyze control flows, we constructed function-level invocation trees from transaction data, aggregating invocations targeting the identified protocol contracts (detailed methodology provided in Section 4). We simplified these extracted control flows by excluding read-only functions, as they do not alter contract state and are inherently benign. Furthermore, we omitted five standardized ERC20 token functions—transfer, transferFrom, approve, increaseAllowance, and decreaseAllowance—since these functions are extensively tested, widely recognized as secure, and typically exhibit straightforward control flows without branching or external interactions. Such standard functions do not usually expose significant security risks targeted by CrossGuard. Instead, vulnerabilities emerge from more complex and interactive control flows (further discussed in Section 5.2). By applying this simplification, we concentrated on critical control flows that significantly impact protocol security, enabling a clearer understanding of the interactions leading to security breaches.
@@
We assessed these configurations using historical transactions from $30$ benchmarks collected in Section 3. To further evaluate CrossGuard under extreme conditions, we applied it to another $3$ widely adopted DeFi protocols—AAVE, Lido, and Uniswap—which serve as fundamental DeFi building blocks. These protocols attract many DeFi developers and feature the most complex and continuously evolving control flows due to their high composability and extensive integrations. To conduct this evaluation, we collected the core smart contracts for these protocols from their official websites (Aave, 2024; Lido DAO, 2024; Uniswap Labs, 2024). Next, we retrieved the most recent 100,000 transactions interacting with these contracts. We then applied CrossGuard to these transactions, measuring its false positive rate (FP%) under real-world extreme conditions. The experimental results are summarized in Table 3.
DeFi Actors Identification(RQ3): To deeply understand the results of CrossGuard and the diversity of control flows in DeFi protocols, we categorize transactions according to their origins and initiators. We focus explicitly on transactions with non-trivial control flows (nCFs), as these represent complex and less predictable interactions, offering deeper insights into protocol dynamics. We identify four primary actor groups capable of introducing unique, non-trivial control flows: 1. Privileged Transactions (P-Tx): Originated by protocol deployers or administrators via privileged functions (e.g., constructors, administrative operations), typically reflecting protocol setup or administrative management activities. 2. Same Protocol (S-Tx): Transactions initiated by other contracts within the same protocol, developed internally to enhance operational coherence and overall functionality. 3. Other DeFi Protocols (O-Tx): Transactions initiated by externally deployed DeFi protocols (commonly labeled on Etherscan), often through open-source collaboration, enriching the broader DeFi ecosystem. 4. External Individuals (E-Tx): Transactions initiated by contracts deployed by external actors such as arbitrageurs, individual traders, or malicious entities (hackers), generally employing closed-source contracts to exploit vulnerabilities or execute complex strategies.
@@ 6.2. RQ2: Effectiveness of CrossGuard
Victim Protocol Block? FPBlock? FPBlock? FPBlock? FPBlock? FPGas OH(%) Block? FPFPFPNot- Bypassable Flash- Loan
bZx2 4.51 4.51 4.5 3.88 3.57 12.38 0.5 0.31 0.12
Warp 0 0 0 0 0 0.04 0 0 0 $\ast$
CheeseBank 2.06 2.06 2.06 0 0 6.93 0 0 0 $\ast$
InverseFi 14.96 14.94 14.91 0.08 0.08 27.46 0.04 0.03 0.03 $\ast$
CreamFi1 4.30 2.28 1.16 1.15 0.39 4.93 0.13 0.07 0.03
CreamFi2 4.32 3.1 2.27 1.12 0.84 16.95 0.23 0.15 0.09 $\ast$
RariCapital1 1.92 1.25 1.25 1.23 1.22 4.91 0.64 0.53 0.41 $\ast$
RariCapital2 1.42 0.33 0.09 0.09 0.02 82.19 0.02 0.02 0.01
XCarnival 0 0 0 0 0 0.17 0 0 0
Harvest 0 0 0 0 0 0.03 0 0 0 $\ast$
ValueDeFi 0 0 0 0 0 0.07 0 0 0 $\ast$
Yearn 2.47 2.47 2.43 0.33 0 3.86 0 0 0 $\ast$
VisorFi 9.56 9.56 9.56 0 0 43.53 0 0 0
UmbrellaNetwork 0 0 0 0 0 0.07 0 0 0 N/A
PickleFi 0.71 0.71 0.71 0.01 0.01 2.08 0.01 0.01 0.01
Eminence 3.21 3.21 3.21 0 0 15.15 0 0 0 $\ast$
Opyn 0.05 0.05 0 0 0 28.74 0 0 0
IndexFi 5.67 5.64 5.64 0 0 36.85 0 0 0
RevestFi 0 0 0 0 0 36.67 0 0 0
DODO 0.07 0.07 0 0 0 0.99 0 0 0
@@
BeanstalkFarms 5.83 5.83 5.83 0.06 0.06 30.9 0.01 0.01 0.01
DoughFina 0 0 0 0 0 0.16 0 0 0
Bedrock_DeFi 15.41 15.41 15.41 0.26 0.06 25.56 0.06 0.06 0.06 N/A
OnyxDAO 4.49 4.48 4.48 0 0 11.69 0 0 0
BlueberryProtocol 0.65 0 0 0 0 0.33 0 0 0 $\ast$
PrismaFi 31.49 31.49 31.46 1.2 0.7 7.47 0.1 0.05 0.05
PikeFinance 0 0 0 0 0 1.28 0 0 0
DFOX 3.40 3.4 3.39 0 0 30.7 0 0 0 $\ast$
UwULend 2.89 2.55 2.55 2.55 2.38 33.47 0.44 0.35 0.19 $\ast$
Summary 29 2.78 29 2.55 29 2.44 28 0.36 28 0.28 15.52 28 0.07 0.05 0.03
AAVE N/A 9.7 N/A 8.53 N/A 8.52 N/A 8.52 N/A 7.32 34.07 N/A 1.10 0.53 0.22 N/A N/A
Lido N/A 13.09 N/A 13.09 N/A 13.07 N/A 7.29 N/A 7.27 52.12 N/A 4.38 2.04 0.43 N/A N/A
Uniswap N/A 1.17 N/A 1.04 N/A 1.00 N/A 0.49 N/A 0.24 3.25 N/A 0.24 0.23 0.06 N/A N/A
Summary N/A 7.99 N/A 7.55 N/A 7.53 N/A 5.43 N/A 4.94 29.81 N/A 1.91 0.93 0.24
The columns “RQ2” in Table 3 present the results for RQ2. Each configuration is evaluated using two key metrics: “Block?” indicates whether the hack was successfully blocked; “FP%” represents the false positive rate for that configuration. Two sets of benchmarks, $30$ hacked protocols and 3 popular protocols, both include a “Summary” row at the bottom, showing the total number of blocked hacks and the average false positive rate per protocol.
Without the ERC20 heuristic, CrossGuard successfully blocks $29$ out of $30$ hacks. The only exception is UmbrellaNetwork, which falls victim to an integer underflow vulnerability (see Table 2). Since integer underflows are local bugs that do not manifest as anomalous control flows, they cannot be mitigated by control flow integrity mechanisms like CrossGuard. When the ERC20 heuristic is enabled, CrossGuard blocks $28$ out of $30$ hacks. The only exception is Bedrock_DeFi, where the attacker exploited a missing input validation vulnerability by invoking a single vulnerable function, without requiring complex control flows. In addition to this core exploit step, the attacker also executed two ERC20 functions—approve and transferFrom—which, however, were not essential to the attack itself. These ERC20 calls could have been executed as separate transactions and did not contribute to the exploit. Since removing them would not prevent the hack from occurring, CrossGuard does not block this transaction. Overall, for $28$ out of $30$ attack transactions, the exploits involve complex control flows that are effectively captured and blocked by CrossGuard, demonstrating its robustness in preventing sophisticated hacks.
Answer to RQ2: CrossGuard is highly effective, especially when combined with rapid manual feedback, resulting in significant reductions in false positives and improved system reliability.
6.3. RQ3: Control Flows Introduced by Different Actors

### v2

@@ 6. Evaluation
Our evaluation aims to answer the following research questions:
@@ 2. RQ 3:
    How do various actors, aside from hackers, introduce new control flows, and what causes false positives in CrossGuard?
@@ 5. RQ 6:
  bZx Lending (14) oracle manipulation 7
28712 39 1389 22 16060 18 7729 14 3534 14
Warp Lending (75) oracle manipulation 17     ✓     416 1 11 0 404 0 0 0 1 1
CheeseBank Lending (18) oracle manipulation 12     ✓     2377 1 98 0 1690 0 543 0 46 1
InverseFi Lending (41) oracle manipulation 12     ✓     121433 60 705 11 51453 16 46452 1 22823 33
CreamFi1 Lending (21) re-entrancy 6     ✓     190797 77 320 21 181676 20 5794 16 3007 33
CreamFi2 Lending (22) oracle manipulation 24     ✓     220417 218 484 31 204481 40 5046 28 10406 141
RariCapital1 Lending (63) read-only re-entrancy 8     ✓     7137 12 54 1 6831 5 208 9 44 5
RariCapital2 Lending (64) re-entrancy 12     ✓     84423 12 393 1 45217 4 27674 2 11139 5
XCarnival Yield Earning (76) logic flaw 6     ✓     875 3 59 0 800 0 0 0 16 3
Harvest Yield Earning (38) oracle manipulation 11     ✓     30997 7 618 5 30338 1 2 0 39 2
ValueDeFi Yield Earning (72) oracle manipulation 9     ✓     356 1 93 0 262 0 0 0 1 1
Yearn Yield Earning (77) logic flaw 7     ✓     133161 14 2860 10 80618 4 44951 1 4732 3
VisorFi Yield Earning (73) re-entrancy 3
85420 2 52 1 26480 1 47232 0 11656 1
PickleFi Yield Earning (56) fake tokens 4     ✓     7511 7 1402 0 4850 5 1231 1 28 1
Eminence DeFi (27) logic flaw 6     ✓     21345 1 24 0 9402 0 9930 0 1989 1
Opyn DeFi (55) logic flaw 4
3921 2 16 1 607 0 1 0 3297 1
IndexFi DeFi (40) logic flaw 6     ✓     70325 5 84 0 14961 4 28182 0 27098 1
RevestFi Yield Earning (65) re-entrancy 5     ✓     2176 4 30 1 2127 3 11 0 8 1
DODO DeFi (25) access control 2
1519 2 2 1 1285 0 195 0 37 1
Punk NFT (61) access control 3     ✓     108 1 11 0 96 0 0 0 1 1
BeanstalkFarms DAO (8) DAO governance attack 8     ✓     58555 12 238 8 43713 1 8860 0 5744 4
DoughFina Lending (26) no input validation 2     ✓     18 2 17 1 0 0 0 0 1 1
Bedrock_DeFi Restaking (10) price miscalculation 3     ✗     3414 2 4 0 2127 0 893 0 390 2
OnyxDAO DAO (54) fake market 8     ✓     157106 3 212 0 97210 0 357 0 59327 3
BlueberryProtocol Yield Earning (11) decimal difference 5     ✓     464 1 113 0 347 0 0 0 4 1
PrismaFi Restaking (60) no input validation 3     ✓     43564 31 123 1 24722 20 4185 0 14534 23
PikeFinance Lending (57) uninitialized proxy 1     ✓     8408 1 15 0 7026 0 0 0 1367 1
GFOX Game Fi (36) access control 2     ✓     12433 1 28 0 9392 0 12 0 3001 1
UwULend Lending (71) oracle manipulation 1     ✓     19681 90 126 0 18296 43 6 3 1253 72
Audius Lending (5) initialization bug 3         6923 3 8 0 6850 1 4 0 61 2
OmniNFT NFT (53) re-entrancy 2         95 4 9 2 72 1 0 0 14 1
MetaSwap DEX (50) logic flaw 3         7027 4 22 1 6641 2 4 0 360 1
Auctus DeFi (4) access control 1         35 1 1 0 33 0 0 0 1 1
BaconProtocol Yield Earning (7) re-entrancy 1         745 3 5 0 680 0 42 1 18 2
MonoXFi DEX (51) logic flaw 4         1037 1 52 0 921 0 0 0 64 1
NowSwap DEX (52) arbitrary call 1         2290 3 11 1 0 0 0 0 2279 2
PopsicleFi Yield Earning (59) logic flaw 6         2648 2 117 0 2499 0 0 0 32 2
                                  Avg. Ratio 100 100 6.11 18.35 70.14 20.01 10.58 5.97 13.17 65.64
@@ 6.1. Methodology
Hacked Protocol Selection: We systematically selected hacked DeFi protocols that experienced significant financial losses (exceeding \$300k) on Ethereum. Protocols were selected from three complementary sources: (1) victim protocols identified by (Chen et al., 2024) between February 14, 2020 and August 1, 2022; (2) hacked protocols reported by (Zhang et al., 2023) within the same period; and (3) protocols compromised between February and July 2024, as documented by DeFiHackLabs (Contributors, 2025). Cross-chain bridge hacks were excluded.
Target Contract Selection and Protocol Filtering: For each selected protocol, we identified all relevant victim protocol contracts by analyzing deployer addresses and labels on Etherscan (30). We then manually examined the control flow of each hack transaction with respect to these contracts. We filtered out hacks involving offchain components and hacks with only trivial control flows w.r.t. the protocol contracts, in line with our study’s scope (see Section 4). After this filtering process, our final dataset consists of $37$ hack incidents: $21$ from source (1), $8$ from source (2), and $8$ from source (3). Table 2 presents our selected benchmarks, including their Protocol Type, hack transaction links(Hack), exploited vulnerability category (Hack Type), and number of affected contracts (#C). The vulnerabilities span diverse attack vectors including oracle manipulation, re-entrancy, DAO governance attack, access control failures, and input validation errors. Our selected protocols represent a comprehensive spectrum of DeFi categories, ensuring broad applicability of our findings.
Transaction History Retrieval: We then retrieved the transaction history of these identified contracts from their deployment until the hack. This complete transaction history, including the hack transaction itself, constitutes the dataset of each victim protocol.  [^7] Although we might miss a few affected protocol contracts due to varying deployers or indirect involvement, having more contracts to protect would only introduce more complexity to the control flows, not reduce it. Thus, our analysis remains valid even with a subset of core contracts. Then we evaluate CrossGuard on this dataset.
@@
We assessed these configurations using historical transactions from $37$ benchmarks collected in Section 2. To further evaluate CrossGuard under extreme conditions, we applied it to another $3$ widely adopted DeFi protocols(AAVE, Lido, and Uniswap) which serve as fundamental DeFi building blocks. These protocols attract many DeFi developers and feature the most complex and continuously evolving control flows due to their high composability and extensive integrations. To conduct this evaluation, we collected the core smart contracts for these protocols from their official websites (Aave, 2024; Lido DAO, 2024; Uniswap Labs, 2024). Next, we retrieved the most recent 100,000 transactions interacting with these contracts. We then applied CrossGuard to these transactions, measuring its false positive rate (FP%) under real-world extreme conditions. The experimental results are summarized in Table 3.
DeFi Actors Identification(RQ3): To deeply understand the results of CrossGuard and the diversity of control flows in DeFi protocols, we categorize transactions according to their origins and initiators. We focus explicitly on transactions with non-trivial control flows (nCFs), as these represent complex and less predictable interactions, offering deeper insights into protocol dynamics. We identify four primary actor groups capable of introducing unique, non-trivial control flows: 1. Privileged Transactions (P-Tx): Originated by protocol deployers or administrators via privileged functions (e.g., constructors, administrative operations), typically reflecting protocol setup or administrative management activities. 2. Same Protocol (S-Tx): Transactions initiated by other contracts within the same protocol, developed internally to enhance operational coherence and overall functionality. 3. Other DeFi Protocols (O-Tx): Transactions initiated by externally deployed DeFi protocols (commonly labeled on Etherscan), often through open-source collaboration, enriching the broader DeFi ecosystem. 4. External Actors (E-Tx): Transactions initiated by contracts deployed by external actors such as arbitrageurs, advanced DeFi users, or malicious entities (hackers), generally employing closed-source contracts to execute complex strategies or exploit vulnerabilities.
@@ 6.2. RQ2: Effectiveness of CrossGuard
Victim Protocol Block? FP% Block? FP% Block? FP% Block? FP% Block? FP% Gas OH(%) Block? FP% FP% FP% Not- Bypassable Flash- Loan
bZx 4.51 4.51 4.5 3.88 3.57 6.29 0.4 0.23 0.06
Warp 0 0 0 0 0 0.03 0 0 0 $\ast$
CheeseBank 2.06 2.06 2.06 0 0 0.05 0 0 0 $\ast$
InverseFi 14.96 14.94 14.91 0.08 0.08 2.17 0.04 0.03 0.03 $\ast$
CreamFi1 4.30 2.28 1.16 1.15 0.39 3.63 0.13 0.07 0.03
CreamFi2 4.32 3.1 2.27 1.12 0.84 9.23 0.21 0.15 0.09 $\ast$
RariCapital1 1.92 1.25 1.25 1.23 1.22 2.14 0.62 0.5 0.36 $\ast$
RariCapital2 1.42 0.33 0.09 0.09 0.02 5.23 0.02 0.02 0.01
XCarnival 0 0 0 0 0 0.15 0 0 0
Harvest 0 0 0 0 0 0.02 0 0 0 $\ast$
ValueDeFi 0 0 0 0 0 0.06 0 0 0 $\ast$
Yearn 2.47 2.47 2.43 0.33 0 1.01 0 0 0 $\ast$
VisorFi 9.56 9.56 9.56 0 0 0.01 0 0 0
PickleFi 0.71 0.71 0.71 0.01 0.01 0.54 0.01 0.01 0.01
Eminence 3.21 3.21 3.21 0 0 0.7 0 0 0 $\ast$
Opyn 0.05 0.05 0 0 0 11.5 0 0 0
IndexFi 5.67 5.64 5.64 0 0 13.48 0 0 0
RevestFi 0 0 0 0 0 0.09 0 0 0
DODO 0.07 0.07 0 0 0 1.58 0 0 0
@@
BeanstalkFarms 5.83 5.83 5.83 0.06 0.06 19.89 0.01 0.01 0.01
DoughFina 0 0 0 0 0 0.09 0 0 0
Bedrock_DeFi 15.41 15.41 15.41 0.26 0.06 10.51 0.06 0.06 0.06 N/A
OnyxDAO 4.5 4.48 4.48 0 0 0.1 0 0 0
BlueberryProtocol 0.65 0 0 0 0 0.29 0 0 0 $\ast$
PrismaFi 31.49 31.49 31.46 1.2 0.7 0.68 0.05 0.02 0.02
PikeFinance 0 0 0 0 0 0.91 0 0 0
GFOX 3.40 3.4 3.39 0 0 0.02 0 0 0 $\ast$
UwULend 2.89 2.55 2.55 2.55 2.38 10.51 0.41 0.33 0.17 $\ast$
Audius 0.03 0.01 0.01 0.01 0.01 0.94 0.01 0.01 0.01
OmniNFT 0 0 0 0 0 2.33 0 0 0
MetaSwap 0.01 0.01 0.01 0.01 0 3.63 0 0 0
Auctus 0 0 0 0 0 0.02 0 0 0 N/A
BaconProtocol 0.27 0.27 0.27 0.27 0.27 0.75 0.27 0.27 0.27
MonoXFi 0 0 0 0 0 3.52 0 0 0
NowSwap 0.31 0.31 0.31 0.04 0.04 18.08 0.04 0.04 0.04
PopsicleFi 0.42 0.42 0.42 0.08 0.08 0.24 0.08 0.04 0.04 $\ast$
Summary 37 3.26 37 3.09 37 3.03 36 0.33 35 0.26 3.53 35 0.06 0.05 0.03 31 26
AAVE N/A 9.69 N/A 8.53 N/A 8.52 N/A 8.52 N/A 7.32 14.13 N/A 1.10 0.53 0.22 N/A N/A
Lido N/A 13.09 N/A 13.09 N/A 13.07 N/A 7.29 N/A 7.27 6.84 N/A 4.38 2.04 0.43 N/A N/A
Uniswap N/A 1.17 N/A 1.04 N/A 1.00 N/A 0.49 N/A 0.24 1.59 N/A 0.24 0.23 0.05 N/A N/A
Summary N/A 7.98 N/A 7.55 N/A 7.53 N/A 5.43 N/A 4.94 7.52 N/A 1.91 0.93 0.23
The columns “RQ2” in Table 3 present the results for RQ2. Each configuration is evaluated using two key metrics: “Block?” indicates whether the hack was successfully blocked; “FP%” represents the false positive rate for that configuration. Two sets of benchmarks, $37$ hacked protocols and $3$ popular protocols, both include a “Summary” row at the bottom, showing the total number of blocked hacks and the average false positive rate per protocol. When the ERC20 and RAW heuristics are enabled, CrossGuard blocks $35$ out of $37$ hacks. The only exceptions are Bedrock_DeFi and Auctus. In the case of Bedrock_DeFi, the attacker exploited a missing input validation vulnerability by invoking a single function; while two ERC20 functions were also called, they were not essential to the core exploit mechanism. Similarly, for Auctus, the attacker exploited an access control vulnerability by repeatedly calling the same function to drain funds. In both instances, these operations could have been executed as independent transactions without relying on the complex control flows that CrossGuard is designed to detect. Consequently, CrossGuard did not block these transactions. Overall, for the $35$ blocked attacks, the exploits involved intricate control flows that were effectively captured by CrossGuard, demonstrating its robustness against sophisticated hacks.
6.3. RQ3: Control Flows Introduced by Different Actors and False Positives